OpenStudy (emilyjones284):

Ill give a medal (: If f(x) = x^2 -1, what is the equation for f^–1(x)?

hartnn (hartnn):

y= x^2-1 replace x and y x = y^2-1 now find y from here... start by adding one to both sides :)

OpenStudy (emilyjones284):

haha thanks...but its just the solving part that i dont understand :/

hartnn (hartnn):

no problem, thats where i gave you a head start ADD 1 to both sides, what u get ?

OpenStudy (emilyjones284):

x+1=y^2?

hartnn (hartnn):

good! now take square root on both sides .. and viola, you are done!

OpenStudy (emilyjones284):

ohh okay so the answer is just +or - the square root of x+1?

hartnn (hartnn):

correct :) it looks like \(\large f^{-1}(x)=\pm \sqrt{x+1}\)

OpenStudy (emilyjones284):

oh i got it thank you (:

hartnn (hartnn):

welcome ^_^